40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 1 of 329  Not logged in ELOG logo
New entries since:Wed Dec 31 16:00:00 1969
ID Date Author Type Category Subject
  16590   Fri Jan 14 18:12:47 2022 AnchalSummaryBHDAS4 placed in ITMY Chamber, OSEMs connected

AS4 was succesfully suspended and trasported to ITMY chamber (40m/16589). We placed it near the door to make it easy to tune the OSEMs. We connected the OSEMs and found that the LL OSEM is not responding. Even though the the OSEMs are completely out right now, there was no signal on this OSEM. This could be an issue in either at the LED driver circuit or the PD circuit in AS4 Sat Amp box, or it could be the OSEM that is bad. We'll investigate further next day. For now, we recorded the full brightness reading for the OSEMs:

  • UL: 32767  -> 16383
  • UR: 29420 -> 14710
  • LR: 30100 -> 15050
  • SD: 29222 -> 14611

Another thing to note is that UL value above is not changing at all. I checked the CDS screen and the the ADC input is overflowing in complete bright position of the OSEM.

  16589   Fri Jan 14 17:33:10 2022 YehonathanUpdateBHDAS4 resurrection

{Yehonathan, Anchal}

Came this morning, the gluing of the magnets was 100% successful. Side blocks, counterweights were assembled. We suspend AS4 and adjust the roll balance and the magnet height (attachments 1,2). OpLev was slightly realigned.

The pitch was balanced. We had to compensate for the pitch shift due to the locking of the counterweights. Once we got good pitch balance, the motion spectrum was taken (attachment 3). Major peaks are at 755mHz, 953mHz, 1040mHz.

Previous peaks were 755mHz, 964mHz, and 1.062Hz so not much has changed. We pushed back the OSEMs, adjusted OSEM plate and locked it tightly. We lock the EQ stops and transfer AS4 to the vacuum chamber in foil. We open the foil inside the chamber. No magnets were broken. Everything seems to be intact. We connect the OSEMs to CDS.

Attachment 1: AS4_roll_balance2.png
AS4_roll_balance2.png
Attachment 2: AS4_magnet_height2.png
AS4_magnet_height2.png
Attachment 3: FreeSwingingSpectra.pdf
FreeSwingingSpectra.pdf
  16588   Fri Jan 14 14:04:51 2022 PacoUpdateElectronicsRFSoC 2x2 board arrived

The Xilinx RFSoC 2x2 board arrived right before the winter break, so this is kind of an overdue elog. I unboxed it, it came with two ~15 cm SMA M-M cables, an SD card preloaded with the ARM processor and a few overlay jupyter notebooks, a two-piece AC/DC adapter (kind of like a laptop charger), and a USB 3.0 cable. I got a 1U box, lid, and assembled a prototype box to hold this board, but this need not be a permanent solution (see Attachment #1). I drilled 4 thru holes on the bottom of the box to hold the board in place. A large component exceeds the 1U height, but is thin enough to clear one of the thin slits at the top (I believe this is a fuse of some sort). Then, I found a brand new front panel, and drilled 4x 13/32 thru holes in the front for SMA F-F connectors.

I powered the board, and quickly accessed its tutorial notebooks, including a spectrum analyzer and signal generators just to quickly check it works normally. The board has 2 fast RFADCs and 2 RFDACs exposed, 14 and 12 bit respectively, running at up to 4 GSps.

Attachment 1: PXL_20220114_211249499.jpg
PXL_20220114_211249499.jpg
  16587   Fri Jan 14 13:46:25 2022 AnchalUpdateBHDPR2 transmission calculation updated

I updated the arm cavity roundtrip losses due to scattering. Yehonathan told me that arm cavity looses 50ppm every roundtrip other than the transmission losses. With the updated arm cavity loss:

  PRFPMI LO Power (mW) Unlocked PRC LO Power (uW) PRC Gain
pre-2010 ITM 8 28 15.2
V6:704 24 113 12

 

Attachment 1: LO_power_vs_PR2_transmission.pdf
LO_power_vs_PR2_transmission.pdf
Attachment 2: PRC_Gain_vs_PR2_transmission.pdf
PRC_Gain_vs_PR2_transmission.pdf
Attachment 3: PR2_Trans_Calc.ipynb.zip
  16586   Fri Jan 14 12:01:21 2022 AnchalUpdateElectronicsBS & ITMY feedthroughs labeled and connected to Sat Amps

I labeled all the newly installed flanges and connected the in-air cables (40m/16530) to appropriate ports. These cables are connected to the CDS system on 1Y1/1Y0 racks through the satellite amplifiers. So all new optics now can be damped as soon as they are placed. We need to make more DB9 plugs for setting "Acquire" mode on the HAM-A coil drivers since our Binary input system is not ready yet. Right now, we only have 2 such plugs which means only one optic and be damped at a time.

 

  16585   Fri Jan 14 11:00:29 2022 AnchalUpdateBHDPR2 transmission calculation

Yeah, I counted the loss from arm cavities as the transmission from ETMs on each bounce. I assumed Michelson to be perfectly aligned to get no light at the dark port.  Should I use some other number for the round-trip loss in the arm cavity?

  16584   Fri Jan 14 03:07:04 2022 KojiUpdateBHDPR2 transmission calculation

I opened the notebook but I was not sure where you have the loss per bounce for the arm cavity.

    PRC_RT_Loss = 2 * PR3_T + 2 * PR2_T + 2 * Arm_Cavity_Finesse * ETM_T + PRM_T

Do you count the arm reflection loss to be only 2 * 13ppm * 450 = 1.17%?

  16583   Thu Jan 13 17:10:55 2022 AnchalUpdateBHDPR2 transmission calculation

I corrected the calculation by adding losses by the arm cavity ends times the arm cavity finesse and also taking into account the folding of the cavity mirror. I used exact formula for finesse calculation and divided it by pi to get the PRC gain from there. Attachment 3 is the notebook for referring to the calculations I made.

Note that using V6-704 would provide 35 mW of LO power when PRFPMI is locked and 113 uW for alignment, but will bring down the PRC Gain to 17.5.

pre-2010 ITM (if it is still an option) would provide 12 mW of LO power when PRFPMI is locked and 28 uW for alignment, but will keep the PRC Gain to 24.6.

I still have to do a curvature check on the V6-704 optic.

Attachment 1: LO_power_vs_PR2_transmission.pdf
LO_power_vs_PR2_transmission.pdf
Attachment 2: PRC_Gain_vs_PR2_transmission.pdf
PRC_Gain_vs_PR2_transmission.pdf
Attachment 3: PR2_Trans_Calc.ipynb.zip
  16582   Thu Jan 13 16:08:00 2022 YehonathanUpdateBHDgluing magnets after AS1/4 misfortune

{Yehonathan, Anchal, Paco}

In the cleanroom, we removed AS1 and AS4 from their SOS towers. We removed the mirrors from the adapters and put them in their boxes. The broken magnets were collected from the towers and their surfaces were cleaned as well as the magnet sockets on the two adapters and on the side block from where the magnets were knocked off.

We prepared our last batch of glue (more glue was ordered three days ago) and glue 2 side magnets and 2 face magnets. We also took the chance and apply glue on the counterweights on the thick optic adapters so there is no need to look for alternatives for now.

The peek screws and nuts were assembled on the thick optics SOS towers instead of the metal screws and nuts that were used as upper back EQ stops.

  16581   Thu Jan 13 12:29:27 2022 AnchalSummaryBHDAS4 LR magnet broke

After the debacle with AS1 (40m/16580), we decided the put the PEEK earthquake stop by first removing the lower OSEM plate and then doing it. So I unfastened AS4 from its position with the earthquake stops in place and moved the suspension to the center of the table. Then I carefully removed the bottom OSEM plate. But I found out that the LR magnet is broken and lying on the floor of the suspension sad. Given my past on the same day, it could be me breaking it again during the moving of the suspension of taking off the OSEM plate or there is a small chance that this break happened before. Regardless of fault, this meant we have to resuspend AS4 again as well. So we transported AS4 back to the clean room and the work on it's re-suspension has begun.

  16580   Thu Jan 13 12:24:08 2022 AnchalSummaryBHDAS1 SD and LR magnets broke

[Anchal (vacuum work), Paco (outside)]

After the AS1 Sat Amp fix (40m/16579), we today were able to tune all OSEMs to the mid-bright level. But when we were about to call it, we were told that the new PEEK earthquake stop screw and bolts need to go on the thin suspended optics. Against better judgment, we decided to install the new back earthquake stop in-situ since we had tuned all OSEMs already. I installed the new stop but after that found that in the process I have broken off the side magnet and LR magnet from the optic adaptor and they are inside the OSEM coils now. This means we'll have to redo the AS1 suspension almost from scratch again sad. We have transported AS1 to the cleanroom where the work on re-suspension has begun.

  16579   Thu Jan 13 09:48:41 2022 AnchalSummaryBHDAS1 Sat Amp fixed

I fixed the issue in AS1 Sat Amp (S2100741) by using a razor blade. I cut the short between the two places, cleaned up the area and covered it with electrical tape. However, later feedback from Rana was to not use electrical tape as it dries up and becomes brittle and lfaky in long run. So after the AS1 OSEM tuning is over, I'll open this box again and use something else to insulate the exposed area. See attached pictures for current status.

 

Attachment 1: signal-2022-01-13-094823_001.jpeg
signal-2022-01-13-094823_001.jpeg
Attachment 2: signal-2022-01-13-094823_002.jpeg
signal-2022-01-13-094823_002.jpeg
  16578   Tue Jan 11 18:40:25 2022 AnchalSummaryBHDAS1 Sat Amp has a PCB issue

AS1 Sat Amp (S2100741) has a critical PCB issue on it's Ch5-8 board S2100548. This board is supposed to just feed through the coil driver signal from the front DB9 connector to the back DB25 connector but it has a short between pins 2 and 7 at the "Coil Input" end (connector J1). The short persists even after I disconnect the sat amp to the flange connector on the back of this board, which definitely means the short is present in the passive channeling through the PCB or at the soldering points of the two DB connectors. I just flipped the board and found that the soldering connections are clean and separate. I think we'll have to use one of the spare sat amp boxes for AS1 for now, while we either declare this one manufacture defected or fix the issue.

I actually found the short on the PCB trace by just looking carefully at it. Attachment 1 shows the photo of it. Maybe we can fix this by simply cutting the tumor between the two traces (why are these traces so close together in such a large board anyways!!!), but I'm not sure if that is a reliable way of fixing this issue. I'll wait for Koji's comments on what to do with this. We'll recommence OSEM tuning for AS1 tomorrow with fixed electronics.

Attachment 1: signal-2022-01-11-184917.jpeg
signal-2022-01-11-184917.jpeg
  16577   Tue Jan 11 18:18:29 2022 AnchalSummaryBHDAttempted OSEM installation on AS1

[Anchal, Paco, Yehonathan]

Connected in-cir cable to new flange on ITMY Chamber

Connected OSEM one-by-one. Starting from top right  to left (PIn1)

1st connector: LL -> UR -> UL

2nd connector: LR -> SD

Loosening all OSEMs and taking them out and noting full bright readings:

  • SD: 29564 -> 14787
  • LR: 30902 -> 15451
  • UR: 29280 -> 14640
  • LL: 27690 -> 13845
  • UL: 27668 -> 13834

:( We had to stop here as we were unable to actuate on the side coils. We checked the signal chain and found that the monitor output of AS1 LL/SD coil driver is responding to offset changes in the coil output filter module of AS1 side. However, when we connected the output of the coil driver through a breakout board to the AS1 Sat Amp, we saw no signal. We tried switching the coil driver bo with another one one the rack but we found the exact same issue. This led us to believe that something must be wrong with the AS1 Sat Amp. We checked the output of the AS1 LL/SD coil driver without connecting it to the sat amp and found that the output was responding to our CDS changes. Then we checked the second "Coil Input" port of the AS1 Sat Amp, and found that pins 2-7 and pins 3-8 are shorted. This means channel 5 and 8 on this box would be shorted. This is the reason why we were unable to actuate on the coils. I'll work on debugging this box, my first guess is that the ribbon cable is bad.

  16575   Tue Jan 11 15:21:16 2022 AnchalUpdateBHDPR2 transmission calculation

I did this simple calculation where I assumed 1W power from laser and 10% transmission past IMC. We would go ahead with V6-704/V6-705 ATFilms 3/8" optic. It would bring down the PRC gain to ~30 but will provide plenty of light for LO beam and alignment.

Attachment 1: LO_power_vs_PR2_transmission.pdf
LO_power_vs_PR2_transmission.pdf
Attachment 2: PRC_Gain_vs_PR2_transmission.pdf
PRC_Gain_vs_PR2_transmission.pdf
Attachment 3: PRS_Trans_Calc.ipynb.zip
  16574   Tue Jan 11 14:21:53 2022 PacoUpdateElectronicsBS feedthroughs and in-vac cables installed

[Paco, Yehonathan, Chub]

The BS chamber 10" flange with 4 DSUB-25 feedthroughs has been installed with the cables connected at the in-vac side. This is the second of two flanges, and includes 4 cables ordered vertically in stacks of 2 & 2 for [[LO2-1, LO2-2, PR3-1, PR3-2]] respectively.

  16573   Tue Jan 11 13:43:14 2022 KojiUpdateSUSTemporary watchdog

I don't remember the syntax of the db file, but here this calc channel computes A&B. And A&B corresponds to INPA and INPB.

        field(CALC,"A&B")
        field(INPA,"C1:SUS-LO1_UL_COMM  PP  NMS")
        field(INPB,"C1:SUS-LO1_LATCH_OFF  PP  MS")

What is this LATCH doing?

 

  16572   Tue Jan 11 12:19:12 2022 AnchalSummaryBHDLO1 Input Matrix Diagonalization performed.

The frree swinging test was successful. I ran the input matrix diagonalization code (scripts/SUS/InMAtCalc/sus_diagonalization.py) on the LO1 free swinging data collected last night. The logfile and results are stroed in scripts/SUS/InMatCalc/LO1 directory. Attachment 1 shows the power spectral density of the DOF bassis data (POS, PIT, YAW, SIDE) before and after the diagonalization. Attachment 2 shows the fitted peaks.


Free Swinging Resonances Peak Fits
  Resonant Frequency [Hz] Q A
POS 0.941 506 84
PIT 1.015 304 778
YAW 0.694 300 626
SIDE 0.999 371 49

LO1 New Input Matrix
  UL UR LR LL SIDE
POS
0.12
0.137
0.338
0.321
0.004
PIT
1.282
1.087
-0.57
-0.375
-0.843
YAW
1.07
-0.921
-1.081
0.91
0.098
SIDE
-0.042
0.383
0.326
-0.099
0.857

The new matrix was loaded on LO1 input matrix and this resulted in no control loop oscillations at least. I'll compare the performance of the loops in future soon.

Attachment 1: LO1_SUS_InpMat_Diagnolization.pdf
LO1_SUS_InpMat_Diagnolization.pdf
Attachment 2: LO1_FreeSwingData_PeakFitting.pdf
LO1_FreeSwingData_PeakFitting.pdf
  16571   Tue Jan 11 10:58:58 2022 TegaUpdateSUSTemporary watchdog

Started working on this. First created a git repo for tracking https://git.ligo.org/40m/susaux.git

I have looked through the folder to see what needs doing and I think I know what needs to be done for the final case by just following the same pattern for the other optics, which I am listing below

- Create database file for the BHD optics, say C1_SUS-AUX_LO1.db by copying another optic database file say C1_SUS-AUX_SRM. Then replace the optic name.

- Insert a new line "C1:SUS-LO1_LATCH_OFF" in the file autoBurt_watchdogs.req

- Populate the file autoBurt.req with the appropriate channels for LO1

- Populate the file C1SUSaux_post.sh with the corresponding commands for LO1

- Add the line dbLoadDatabase("/cvs/cds/caltech/target/c1susaux/C1_SUS-AUX_LO1.db") to the file C1SUSaux.cmd

 

For the temporary watchdog, we comment everything I have just talked about, and do only what come next.

My question is the following:

I understand that we need to use the OUT16 slow channel as a temporary watchdog since we don't currently have access to the slow channels bcos the Acromag units have not been installed. My guess from Koji's instructions is that we need to update the channels in the last two fields "INPA" and "INPB" below

record(calc,"C1:SUS-LO1_UL_CALC")
{
        field(DESC,"ANDs Enable with Watchdog")
        field(SCAN,"1 second")
        field(PHAS,"1")
        field(PREC,"2")
        field(HOPR,"40")
        field(LOPR,"-40")
        field(CALC,"A&B")
        field(INPA,"C1:SUS-LO1_UL_COMM  PP  NMS")
        field(INPB,"C1:SUS-LO1_LATCH_OFF  PP  MS")
}

Suppose we replace the channel for INPA with C1:SUS-LO1_ULCOIL_OUT16, what about INPB. Is this even the right thing to do as I am just guessing here?

 

  16570   Tue Jan 11 10:46:07 2022 TegaUpdateCDSSUS screen debugging

Seen. Thanks.

Red Arrow: The channel was labeled incorrectly as INMON instead of OUTPUT

Green Arrow: OK, I will create a custom medm screen for this.

Blue arrow: Hmm, OK I will look into this. Doing this work remotely is a pain as the medm response is quite slow for poking around.

Orange circle: OK, I'll move this to the left side of the line.

Note to self: I also noticed another error on the side (LPYS blue box just b4 the sum). The channel is pointing to YAW instead of the side, so needs to be fixed as well.

Quote:

Indicated by the red arrow:
Even when the side damping servo is off, the number appears at the input of the output matrix

Indicated by the green arrows:
The face magnets and the side magnets use different ADCs. How about opening a custom ADC panel that accommodates all ADCs at once? Same for the DAC.

Indicated by the blue arrows:
This button opens a custom FM window. When the pitch gain was modified with a ramping time, the pitch and yaw gain grows at the same time even though only the pitch gain was modified.

Indicated by the orange circle:
The numbers are not indicated here, but they are input-related numbers (for watchdogging) rather than output-related numbers. It is confusing to place them here.

 

  16569   Tue Jan 11 10:23:18 2022 PacoUpdateElectronicsITMY feedthroughs and in-vac cables installed - part II

[Paco, Chub]

The ITMY 10" flange with 4 DSUB-25 feedthroughs has been installed with the cables connected at the in-vac side. This is the second of two flanges, and includes 4 cables ordered vertically in stacks of 2 & 2 for [[AS1-1, AS1-2, AS4-1, AS4-2]] respectively. No major incidents during this one, except maybe a note that all the bolts were extremely dirty and covered with gunk, so we gave a quick swipe with wet cloths before reinstalling them.

  16568   Tue Jan 11 09:53:14 2022 not KojiUpdateBHDSOS assembly -- Peek screws and nuts

I handed the Peek parts we got from McMaster to Jordan for C&B.

  16567   Mon Jan 10 18:36:41 2022 AnchalSummaryBHDLO1 free swinging test set to trigger

LO1 is set to go through a free swinging test at 1 am tonight. We have used this script (scripts/SUS/InMatCalc/freeSwing.py) reliably in the past so we expect no issues, it has a error catching block to restore all changes at the end of the test or if something goes wrong.

To access the test, on rossa, type:

tmux a -t freeSwingLO1

Then you can kill the script if required by Ctrl-C, it will restore all changes while exiting.

  16566   Mon Jan 10 18:20:45 2022 AnchalUpdateBHDTested 2" PR2 candidates transmission

I tested 2 more optics found by Paco and Yehonathan in QIL.

  Polarization Incident Power [mW] Transmitted Power [mW] Transmission [ppm]
V6-704 V6-706 p-pol 850 17.1 20118
Yellow cylindrical box p-pol 850 <1 ( could not even see it to measure it with a more sensitive power meter) <1000

I would like someone to redo the second test. I'm not sure what was happening but I could not find the transmitted beam at all on my card even with all lights out. This is either too good a coating and not useful for us or I did something wrong while measuring it.

V6-704, V6-706 mirror seemed like a good candidate as the paper with it said it would be a 200 ppm mirror. But I measured a lot more transmission than that. Now that I see that paper more carefully, it is a 45 degree s-pol mirror, probably that's why it had so much transmission for p-pol at near-normal incidence.

 

  16565   Mon Jan 10 17:04:47 2022 AnchalUpdateBHDAS1 Sat Amp CH2 had offset

We found that there was a small offset (~300 mV) at TP6 and TP8, in PD2 circuit (CH2 of the board). I replaced U3 AD822ARZ but did not see any affect. I disconnected the adaptor board in the back and saw that the offset went away. This might mean that the cable had some flaky short to a power supply pin. However, when I just reinserted the adaptor board back again, there was no offset. We could not find any issue with the board after that to fix, so we left it as it is. If this board gives offset issues in the future, most probably the ribbon cable would be the suspect.

Now all ADC channels are showing no offset or overflows in C1SU2 chassis.

  16564   Mon Jan 10 15:59:46 2022 KojiUpdateBHDPR2 Sat Amp has a bad channel

The issue was present in the cable between the small adapter board and the rear panel. The cable and the Dsub 25 connectors were replaced. The removed parts were resoldered. Did the basic test of the channel.


Attachment 1: I cleaned up the area of the PD3 circuit of S2100556 and checked the voltage when the circuit was energized. The PD photocurrent line from the rear panel had S2100556 even with R25 removed. So the problem was between the rear panel to the outer side of R25. I've started to remove the cables to localize the issue and found that the issue disappeared when the ribbon cable was removed.

Attachment 2: I didn't investigate how the ribbon cable was bad. It was just trashed. The cable and the 25pin Dsub connectors were replaced and the line in question looked normal.

Attachment 3: All the components removed were stuffed again. The I/V-output of the circuit showed a 0.7mV offset but it seemed within the normal range. By touching R25 with a finger made it up to ~10mV as the other channels do. BTW: For 1000pF cap (C10) I used a stock 1000pF cap (KEMET, C330C102JDG5TA, 5%, 1kV, C0G) instead of nominal one (KEMET, C317C102G1G5TA,  2%, 100V, C0G).

Attachment 4: Noticed that the jumpers for shield grounding were missing. So they were installed (Attachment 5). This jumper is connected to Pin13. This line becomes Pin1 of the Dsub25 sat-amp cable because of the adapter board D2100148. The sat amp cable is D2100675. Hmm. In fact, this line does not touch the shield anywhere (unlike the aLIGO case). So only the chassis provides the cable shielding, no matter how the jumpers are connected or not connected.

Attachment 6: Final state of the circuit

Attachment 1: trouble_shoot1.jpg
trouble_shoot1.jpg
Attachment 2: trouble_shoot2.jpg
trouble_shoot2.jpg
Attachment 3: S2100556_PD3.jpg
S2100556_PD3.jpg
Attachment 4: shield_grounding_before.jpg
shield_grounding_before.jpg
Attachment 5: shield_grounding_after.jpg
shield_grounding_after.jpg
Attachment 6: S2100737.jpg
S2100737.jpg
  16563   Mon Jan 10 15:45:55 2022 PacoUpdateElectronicsITMY feedthroughs and in-vac cables installed - part I

The ITMY 10" flange with 10 DSUB-25 feedthroughs has been installed with the cables connected at the in-vac side.  This is the first of two flanges, and includes 5 cables ordered vertically in stacks of 3 & 2 for [[OMC-DCPDs, OMC-QPDs, OMC-PZTs/Pico]] and [[SRM1, SRM2]] respectively from right to left. During installation, two 12-point silver plated bolts were stripped, so Chub had to replace them.

  16562   Mon Jan 10 14:52:51 2022 AnchalSummaryBHDLO1 OSEMs roughly calibrated and noise measured

I used the open light level output of 908 for ITMX side OSEM from 40m/16549 to roughly calibrate cts2um filter module in LO1 OSEM input filters. All values were close to 0.033. As the calibration reduces the signal value by about 30 times, I increased all damping gains by a factor of 30. None of loops went into any unstable oscillations and I witnessed damping of kicks to the optic.


In-loop power spectrum

I also compared in-loop power spectrum of ETMX and LO1 while damping. ETMX was chosen because it is one of the unaffected optics by the upgrade work. ITMX is held by earthquake stops to avoid unnecessary hits to it while doing chamber work.

Attachment 1 and 2 show the power spectrum of in-loop OSEM values (calibrated in um). At high frequencies, we see about 6 times less noise in LO1 OSEM channel noise floor in comparison to ETMX. Some peaks at 660 Hz and 880 Hz are also missing. At low frequencies, the performance of LO1 is mostly similar to EMTX except for a peak (might be loop instability oscillation) at 1.9 Hz and another one at 5.6 Hz. I'll not get into noise hunting or loop optimization at this stage for the suspension. For now, I believe the new electronics are damping the suspensions as good as the old electronics.

Attachment 1: LO1_vs_ETMX_OSEM_Spectrum_LF_x30_Gain.pdf
LO1_vs_ETMX_OSEM_Spectrum_LF_x30_Gain.pdf
Attachment 2: LO1_vs_ETMX_OSEM_Spectrum_HF_x30_Gain.pdf
LO1_vs_ETMX_OSEM_Spectrum_HF_x30_Gain.pdf
  16561   Mon Jan 10 14:00:44 2022 not KojiUpdateBHDSOS assembly -- SR2

Yes,

For the thin optics adapter design, we want Peek 1/4-20 screw (part # 98885A131) to replace the lower back long EQ stop. On it, we will have a Peek washer (part # 93785A600) fastened between two Peek nuts (part #98886A813).

For the thick optics adapter design, we want Peek 1/4-20 screw (part # 98885A131) to replace both the upper and lower back EQ stop. On the upper stop, we need a single Peek nut (part #98886A813).

I will cure-test the Vacseal.

Quote:

Vacseal in the freezer. It could have been expired sooooo many years ago, We need some cure testing.

Can you release the part numbers of the ordered components (and how/where to use them), so that we can incorporate them into the CAD model?

Quote:

Again, we should apply some glue to the counterweight to prevent it from spinning on the setscrew. Is there a glue other than EP30 that we can use?

Related: Peek nuts, screws and washers were ordered from Mcmaster.

 

  16560   Mon Jan 10 13:35:52 2022 AnchalUpdateBHDPR2 Sat Amp has a bad channel

The unit was tested before by Tege. The test included testing the testpoint voltages only. He summarized his work in this doc. The board number is S2100737. Here are the two comments about it:
"This unit presented with an issue on the PD1 circuit of channel 1-4 PCB where the voltage reading on TP6, TP7 and TP8 are -15.1V,  -14.2V, and +14.7V respectively, instead of ~0V.  The unit also has an issue on the PD2 circuit of channel 1-4 PCB because the voltage reading on TP7 and TP8 are  -14.2V, and +14.25V respectively, instead of ~0V."

"Debugging showed that the opamp, AD822ARZ, for PD2 circuit was not working as expected so we replaced with a spare and this fixed the problem. Somehow, the PD1 circuit no longer presents any issues, so everything is now fine with the unit."

Note:  No issues were reported on PD3 circuit is is malfunctioning now.

Quote:

Also: Was this unit tested before? If so, what was the testing result at the time?

 

  16559   Sat Jan 8 16:01:42 2022 PacoSummaryBHDPart IX of BHR upgrade - Placed LO2 filters

Added input filters, input matrix, damping filters, output matrix, coil filters, and copy the state over from ITMX into LO2 screen in anticipation for damping.

  16558   Fri Jan 7 18:28:13 2022 KojiUpdateBHDPR2 Sat Amp has a bad channel

Leave the unit to me. I can look it at on Mon. For a while, you can take a replacement unit from the electronics stack.

Also: Was this unit tested before? If so, what was the testing result at the time?

  16557   Fri Jan 7 18:24:25 2022 KojiUpdateBHDSOS assembly -- SR2

Vacseal in the freezer. It could have been expired sooooo many years ago, We need some cure testing.

Can you release the part numbers of the ordered components (and how/where to use them), so that we can incorporate them into the CAD model?

Quote:

Again, we should apply some glue to the counterweight to prevent it from spinning on the setscrew. Is there a glue other than EP30 that we can use?

Related: Peek nuts, screws and washers were ordered from Mcmaster.

  16556   Fri Jan 7 17:59:45 2022 YehonathanUpdateBHDSOS assembly -- SR2

{Yehonathan, Paco}

{Paco, Yehonathan}

Today we suspended SR2 (E1800089) which Anchal has loaded into the thick optic adapter. Attachments 1,2 show the height and roll balance adjustments.

I realigned the opLev setup and balanced the suspended mass. Attachment 3 shows the motion spectra on the QPD. There are major peaks at 723 mHz, 832 mHz, and 996 mHz. I inserted OSEMs and tightened them in place. I adjusted the OSEM plates to make sure the magnets are at the center of the OSEMs, then I tightened the OSEM plates to the SOS tower.

The optic was locked keeping the alignment fixed on the center of the QPD.

Again, we should apply some glue to the counterweight to prevent it from spinning on the setscrew. Is there a glue other than EP30 that we can use?

Related: Peek nuts, screws and washers were ordered from Mcmaster.

Attachment 1: SR2_roll_balance.png
SR2_roll_balance.png
Attachment 2: SR2_magnet_height.png
SR2_magnet_height.png
Attachment 3: FreeSwingingSpectra.pdf
FreeSwingingSpectra.pdf
  16555   Fri Jan 7 17:54:13 2022 AnchalUpdateBHDPR2 Sat Amp has a bad channel

[Anchal, Paco]

Yesterday we noticed that one of the ADC channels was overflowing. I checked the signal chain and found that CH3 on PR2 Sat Amp was railing. After a lot of debugging, our conclusion is that possible the PD current input trace is shorted to the positive supply through a finite resistance on the PCB. This would mean this PCB has a manufacturing defect. The reason we come to this conclusion is that even after removing the opamp U3 (AD822ARZ), we still measure 12.5 V at the pins of R25 (100 Ohm input resistance)

Please see the schematic for reference. We also checked the resistance between input of R25 (marked PDA above) and positive voltage rail and it came out as 3 kOhms. While I all other channels, this value was 150 kOhms.

I would like it if someone else also takes a look at this. We probably would need to change the PCB in this chassis or use a spare chassis.

  16554   Fri Jan 7 16:17:42 2022 AnchalSummaryBHDPart IX of BHR upgrade - Placed AS1 and AS4 filters

[paco]

Added input filters, input matrix, damping filters, output matrix, coil filters, and copy the state over from LO1 into AS1 screen in anticipation for damping.

Added input filters, input matrix, damping filters, output matrix, coil filters, and copy the state over from LO1 into AS4 screen in anticipation for damping.

  16553   Thu Jan 6 22:18:47 2022 KojiUpdateCDSSUS screen debugging

Indicated by the red arrow:
Even when the side damping servo is off, the number appears at the input of the output matrix

Indicated by the green arrows:
The face magnets and the side magnets use different ADCs. How about opening a custom ADC panel that accommodates all ADCs at once? Same for the DAC.

Indicated by the blue arrows:
This button opens a custom FM window. When the pitch gain was modified with a ramping time, the pitch and yaw gain grows at the same time even though only the pitch gain was modified.

Indicated by the orange circle:
The numbers are not indicated here, but they are input-related numbers (for watchdogging) rather than output-related numbers. It is confusing to place them here.

Attachment 1: Screen_Shot_2022-01-06_at_18.03.24.png
Screen_Shot_2022-01-06_at_18.03.24.png
  16552   Thu Jan 6 21:04:41 2022 AnchalSummaryBHDPart VIII of BHR upgrade - LO1 OSEMs inserted

[Anchal, Koji] Part of elog: 40m/16549.

The magnets on the mirror face are arranged in a manner that the overall magnetic dipole moment is nullified faraway. Because of this, the coil output gains in all such optics need to have positive and negative signs in a butterfly mode pattern (eg. UL, LR: +ve and UR, LL: -ve).

In the particular case of LO1, we chose following coil output gains:

  COIL_GAIN
UL -1
UR 1
LR -1
LL 1
SD -1

This ensures that all damping gains have positive signs. Following damping gain values were chosen:

DOF C1:SUS-LO1_SUSXXX_GAIN
POS 5
PIT 2
YAW 0.2
SIDE 10

Having said that, this is a convention and we need to discuss more on what we want to set a convention (or follow a previous one if it exists). My discussion with Koji came up with the idea of fixing the motion response of an OSEM with respect to coil offset by balancing the coil gains across all optics and use same servo gains for all optics afterwards. But it is a complicated thought coming out of tired minds, needs more discussion.


Important notes for suspending the optics:

  • Do not insert the OSEMs fully. Leave all of the magnet out of the OSEMs before transportation.
  • Tighten the OSEMs completely while adjusting the height of the optic. Adjust height of OSEM holder plate if necessary.
  • Ensure the all cage screws are screwed tight completely.

Photos: https://photos.app.goo.gl/CJsS18vFwjo73Tzs5

  16551   Thu Jan 6 17:16:51 2022 YehonathanUpdateBHDUsing Peek screws/nuts

There were several cases where the long EQ stops didn't perform as expected.

In one type of case, we used a counterweight at the front of the adapter but not in the back leaving a recess where the lower back EQ stop should touch.

In the other type, a recess in the thick optics adapter prevented the upper EQ stop from touching the adapter. In the first thick optic, the screw was screw barely scratched the recess' corner. In the second case, it didn't touch it at all.

In the last group meeting, we discussed using Peek screws (made out of plastic) to prevent metal on metal bumping when the EQ can touch the adapter and Peek nuts when it doesn't to increase its impact area.

Mcmaster has 1.5" long 1/4-20 screws (part number 98885A131) that will fit well in the OSEM plates. We can order 20 of those.

The biggest Peek nuts on Mcmaster however are not big enough (7/16" wide) to cover the entire bottom recess area which is 0.5" wide (they are good enough for the top recess area in the thick adapter optic design). Koji suggested that we can use a big Peek washer for that purpose that can be held between nuts. We should then order 10 Peek nuts (98886A813) and 1 package of 10 Peek washers (0.63" OD) (93785A600).

  16550   Thu Jan 6 17:00:20 2022 YehonathanUpdateBHDSOS assembly -- LO2

{Paco, Yehonathan}

Today we suspended LO2 (E1800089) which Anchal has loaded into the thick optic adapter. Attachments 1,2 show the height and roll balance adjustments.

I realigned the opLev setup and balanced the suspended mass. We figured that if we use 2 counterweights we will be 1 short. We decided to use 1 mass at the back of the adapter. This has the additional advantage that the Viton tip on lower back EQ stop can touch it and act normally. The optic was successfully balanced in this way. Attachment 3 shows the motion spectra on the QPD. There are major peaks at 712 mHz, 854 mHz, 876 mHz, and 996 mHz. As expected using only 1 counterweight raised the center of mass and lowered the pitch resonance frequency. The optic was locked keeping the alignment fixed on the center of the QPD, OSEMs were inserted and the SOS tower was engraved.

We should apply some glue to the counterweight to prevent it from spinning on the setscrew.

Attachment 1: LO2_roll_balance.png
LO2_roll_balance.png
Attachment 2: LO2_magnet_height.png
LO2_magnet_height.png
Attachment 3: FreeSwingingSpectra.pdf
FreeSwingingSpectra.pdf
  16549   Thu Jan 6 15:10:38 2022 KojiUpdateSUSITMX Chamber work

[Anchal, Koji]

=== Summary ===
- ITMX SD OSEM migration done
- LO1 OSEM insertion and precise adjustment (part 1) done
- LO1 POS/PIT/YAW/SD motions were damped


=== General Remarks ===
- 15:00 Entered into ITMX.
- We were equipped with N95 and took physical distance as much as possible.
- 17:00 Temporarily came out from the lab.
- 18:30? Came into the chamber again
- 20:00 Sus damped. OSEM work continues
- 21:00 OSEM installation work done. Exit.

=== ITMX SD OSEM position swap ===
- Moved the LO1 suspension to the center of the chamber
- Removed the ITMX SD OSEM from the right side (west side) and tried to move it to the other side.
- Noted that the open light output of the ITMX SD was 908 at the output of the SDSEN filter module. So the half-light target is 454. These numbers include the "cnt2um" calibration of 0.36. That means the open light raw ADC count was supposed to be 2522.

- The OSEM set screw (silver plated, with a plunger) was removed from the old position. We first tried to recycle it to the other side, but it didn't go into the thread with fingers. After making ourselves convinced that the threaded hole was identical for both sides, we decided to put the new identical plunger set screw with an Allen-key was used to put it in and it went in!
- Now the ITMX SD OSEM was inserted from the east side. Once we saw some shadow on the OSEM signal, the SD damping was turned on with the previous setting. And this successfully damped the side motion. ⭕️
- A bit finer adjustment has been done. After a few trials, we reached the stable output of ~400. Considering the temporary leveling of the table, we decided this is enough for now ⭕️. The set screw was tightened.
- To make the further work safer w.r.t the ITMX magnets, Anchal fastened the EQ stops of the ITMX sus except for the bottom four.
- Photo: [Attachment 1]

=== LO1 OSEM installation ~ wiring ===
- Now LO1 was moved back to the planned position.
- For the wiring, we (temporarily) clamped the in-vac DSUB cables to the stack table with metal clamps.
- Started plugging the OSEMs into the DSUB cables.
- Looking at the LO1-1 cable from the mating side with the longer side top: The top-right pin of the female connector is Pin1 as usual. From right to left LL / UR / UL coils were inserted one by one while looking at the OSEM PD signals.
- LO1-2 cable has the LR / SD coils (from the right to the left) were connected.
- Photo: [Attachment 2]

- LO1 Open light levels (raw ADC counts) the 2nd number is the target half-light level

  • UL 27679 (-> 13840)
  • UR 29395 (-> 14697)
  • LR 30514 (-> 15257)
  • LL 27996 (-> 13998)
  • SD 26034 (-> 13017)

=== RTS Filter implementation ===

- Anchal copied the filter module settings from other suspensions.
- We also implemented the simple input and output matrices.

=== LO1 OSEM insertion ===

- We struggled to make the suspension freely swinging with the OSEMs inserted.

- It seemed that the magnets were sucked to the OSEMs due to magnetic components.
- It turned out that the OSEMs were not fastened well and not seated in the holder plates.
- Once this was fixeded, we found that the mirror height is too high for the given OSEM heights.
  The suspension height (or the OSEM height should be decided with the OSEMs not inserted but fully fastened to prevent misalignment of them.

- Decided to lift up the OSEM plates in situ.
- Soon we found that the OSEM holder plates are not fastened at all [Attachment 3 arrows]
- The plates were successfully lifted up and
the suspension became much more freely swinging even with the OSEMs inserted. ⭕️

=== LO1 damping and more precise OSEM insertion ===

- Once the OSEMs were inserted to the light level of 30~70%, we started to try to dampen the motion. The side damping was somewhat successful, but the face ones were not.
- We checked the filters and found the coil output filters didn't have the alternating signs.
- Once the coil signs were corrected, the damping became more straight forward.
- And the robust damping allowed us the fine-tuning of the OSEM insertion.

- In the end, what we had for the light levels were

  • UL 14379 (52%)
  • UR 14214 (48%)
  • LR 14212 (47%)
  • LL 12869 (46%)
  • SD 14358 (55%)

The damping is working well. [Attachment 4]


Post continues at 40m/16552.

Attachment 1: PXL_20220107_044739280.MP.jpg
PXL_20220107_044739280.MP.jpg
Attachment 2: PXL_20220107_044958224.jpg
PXL_20220107_044958224.jpg
Attachment 3: PXL_20220107_044805503.NIGHT.jpg
PXL_20220107_044805503.NIGHT.jpg
Attachment 4: Screen_Shot_2022-01-06_at_20.54.04.png
Screen_Shot_2022-01-06_at_20.54.04.png
  16548   Thu Jan 6 14:08:14 2022 KojiUpdateCDSMore BHD SUS screens added to sitemap

More BHD SUS screens added to sitemap (Attachment 1)

Attachment 1: Screenshot_2022-01-06_14-06-15.png
Screenshot_2022-01-06_14-06-15.png
  16547   Thu Jan 6 13:54:28 2022 KojiUpdateCDSYearly DAQD fix 2022!

Just restarting all the c1sus2 models fixed the issue. (Attachment 1)

SUS2 ADC1 CH21 is saturated. I'm not yet sure if this is the electronics issue or the ADC issue.
SUS2 ADC1 CH10 also has large offset. This should also be investiagted.

Attachment 1: Screenshot_2022-01-06_13-57-40.png
Screenshot_2022-01-06_13-57-40.png
  16546   Thu Jan 6 12:52:49 2022 AnchalUpdateCDSYearly DAQD fix 2022!

Just as predicted, all realtime models reported "0x4000" error. Read the parent post for more details. I fixed this by following the instructions. I add folowing lines to the file /opt/rtcds/rtscore/release/src/include/drv/spectracomGPS.c in fb1:

/* 2020 had 366 days and no leap second */
       pHardware->gpsOffset += 31622400;
/* 2021 had no leap seconds or leap days, so adjust for that */
       pHardware->gpsOffset += 31536000;

Then is made the package and reloaded it after stoping the daqd services. This brought back all the fast models except C1SUS2 models which are in red due to some other reason that I'll investigate further.

 

  16545   Thu Jan 6 11:54:20 2022 AnchalSummaryBHDPart IX of BHR upgrade - Placed AS1 and AS4

[Paco (Vacuum Work), Anchal]

Today we opened the ITMY Chamber and installed suspended AS1 and AS4 in their planned positions. In doing so, we removed the razor or plate mounted on a pico motor at the south end of the table (see 40m/16450). We needed to make way for AS4 to be installed.


Photos: https://photos.app.goo.gl/YP2ZZhQ3jip3Uhp5A


We need more dog clamps for installing the suspensions, we have used temporary clamps for now. However, knows where new C&B clamps are, please let us know.

  16544   Wed Jan 5 19:18:06 2022 YehonathanUpdateBHDSOS assembly -- AS4

{Paco, Yehonathan, Anchal}

Today we suspended AS4 (E2000226-B). Anchal mounted Lambda Optic mirror with an RoC closest to AS4 in a thin optic mount. He noted that this optic as well as AS1 don't have a wedge angle. The specs claim that the wedge angle is 2 degrees what should have been clearly seen by inspecting the optic with a naked eye. All the ghost beam deflections probably come from the curvature of the mirror.

We did all the height and roll balancing using a camera (Attachment 1,2). We balanced that pitch of the adapter using a QPD not before we realigned the OpLev setup.

We measured the motion spectra (attachment 3). Major peaks are found at 755 mHz, 964 mHz, and 1.062Hz. I locked the counterweights setscrew and observed that the pitch balance doesn't change. I locked the EQ stops such that the alignment of the mirror remained the same by monitoring the QPD signals. I clamped the suspensions wires to the suspension block.

The only thing remaining is inserting the OSEMs.

 

Attachment 1: AS4_roll_balance.png
AS4_roll_balance.png
Attachment 2: AS_4_magnet_height.png
AS_4_magnet_height.png
Attachment 3: FreeSwingingSpectra.pdf
FreeSwingingSpectra.pdf
  16543   Wed Jan 5 17:46:04 2022 AnchalUpdateBHDTested 2" PR2 candidates transmission

I tested 2 more optics today, the old PR2 that we took out and another optic I found in QIL. Both these optics are also not good for our purpose.

 

Polarization Incident Power [mW] Transmitted Power [mW] Transmission [ppm]
Existing PR2 p-pol 910 0.004 4.4
V2-1698 & V2-1700 p-pol 910 595 653846

I'll find thw Y1S optic and test that too. We should start looking for alternate solutions as well.

 

  16542   Tue Jan 4 18:27:23 2022 PacoUpdateBHDSOS assembly -- PR3

[yehonathan, paco, anchal]

We continue suspending PR3 today. Yehonathan and Paco suspended the thick optic in its adapter. After fixing some nominal height and undoing any residual roll angle (see Attachments 1,2 for pictures), we noticed a problem with the pitch angle, so we insert the counterweights all the way in. Nevertheless, we soon found out that we needed to shift one of the two counterweights to the back of the adapter side (so one on each side) in order to tare the pitch angle. This is a newly experienced maneuver that may apply for further thick optics.

After taring the pitch angle roughly, we noted another issue. The wedge (~ 1 deg) on the optic made it such that the protruding socket heads on the thick side bumped against the lower clamp (not the earthquake stop tip itself). Attachments #4,5 show the before/after situation which was solved provisionally by replacing the socket head screws with lower profile (flat) head screws in situ. Again, this operation was highly delicate and specific to wedged thick optics, so for future SOS we should keep it in mind.

Another issue that we had with the new thick optic adapters is that for some reason there is a recession in the upper backside of the adapter (attachment coming soon). This makes the upper back EQ stop too short to touch the adapter. We replaced it with a longer screw. When inserted it doesn't really hit the back of the adapter. Rather, it touches the corner of the recession, stoping the optic with friction.

While all this was happening, Anchal started mounting AS4 on its adapter. After one of the magnets broke off, he switched to another one and succeeded. This is the next target for suspension. We still need to check the orientation of the wedge. Furthermore, we started a gluing session in the afternoon to prepare as much as possible for further SOS during the week. 3 side magnets were glued to side blocks. 3 magnets were glued to 3 adapters that were missing 1 magnet each.

In the afternoon, Yehonathan and Paco set up the QPD and did all the usual balancing, and then Anchal took the data of which the result is shown in Attachment #3. The major peaks are located at 723mHz, 953mHz, and 1.05Hz. Very similar to the case of the thin optic adapters.

Anchal progressed with OSEM installation, and engraving and yehonathan glued the counterweight setscrew in place. After securing the EQ stops, and wrapping the wires in foil, we declare PR3 is ready to be installed.

Attachment 1: PR3_roll_balance.png
PR3_roll_balance.png
Attachment 2: PR3_magnet_height.png
PR3_magnet_height.png
Attachment 3: FreeSwingingSpectra.pdf
FreeSwingingSpectra.pdf
Attachment 4: PXL_20220104_231742123.jpg
PXL_20220104_231742123.jpg
Attachment 5: PXL_20220104_232809203.jpg
PXL_20220104_232809203.jpg
  16541   Tue Jan 4 18:26:59 2022 AnchalUpdateBHDTested 2" PR2 candidates transmission

I used the rejected light from the PBS after the motorized half-wave plate between PMC and IMC injection path (used for input power control to IMC) to measure the transmission of PR2 candidates. These candidates were picked from QIL (QIL/2696). Unfortunately, I don't think either of these mirrors can be used for PR2.

  Polarization Incident Power [mW] Transmitted Power [mW] Transmission [ppm]
V2-2239 & V2-2242 s-pol 940 0.015 16.0
V2-2239 & V2-2242 p-pol 935 0.015 16.0
V6-704 & V6-705 p-pol 925 21 22703

If I remember correctly, we are looking for a 2" flat mirror with a transmission of the order of 1000 ppm. The current PR2 is supposed to have less than 100 ppm transmission which would not leave enough light for LO path.

I've kept the transmission testing setup intact on the PSL table, I'll test existing PR2 and another optic (which is 0.5" thick unfortunately) tomorrow.

  16540   Mon Jan 3 16:46:41 2022 PacoUpdateBHD1Y1 rack work for SR2, PR2, PR3

[Paco, Anchal]

Continued working on 1Y1 rack. Populated the 6 coil drivers, made all connections between sat amp, AA chassis, DAC, and ADC adapters for SR2, PR2, and PR3 suspensions. Powered all boxes and labeled them and cables where needed. Near the end, we had to increase the current limit on the positive rail sorensen (+18 V) from ~ 7 to > 8.0 Amps to feed all the instruments. We also increased the negative (-18 V) current limit proportionally.

We think we are ready for all the new SOS on this side electronics-wise.


Photos: https://photos.app.goo.gl/GviuqLQviSPo1M3G6

ELOG V3.1.3-